OpenStudy (anonymous):

skech the direction field of the differential equation dx/dt =-2t. Any idea please?

OpenStudy (anonymous):

First, draw your axes. Then, find the slope of a potential curve by plugging in various values of \(t\). For example, when \(t=2\), the slope of the curve \(x(t)\) is given by \(\dfrac{dx}{dt}\bigg|_{t=2}=-2(2)=-4\). Since the slope depends only on \(t\), the slope at any point \((2,x(2))\) is -4.  The drawing's far from perfect, but you get the idea.

OpenStudy (anonymous):

Now try a different point, maybe \(t=0\). The slope is \(\dfrac{dx}{dt}\bigg|_{t=0}=-2(0)=0\), so along the \(x(t)\)-axis, every solution curve will have a horizontal tangent line: